# State Your Intention to Steer Your Attention: An AI Assistant for Intentional Digital Living

### v2

@@ 7.2. Impact of INA on Intended Activity and User Focus
As shown in Figure 6, for the two applications that collect user intentions, INA not only reduces off-task time during task execution but also increases users’ reported alignment between their digital activity and their stated intention. The LLM-estimated off-task ratio [^6] Our technical evaluation in Section 5 provides evidence of its accuracy. is significantly lower with INA than with simple reminder (0.104 vs. 0.166, $p<.001$). Users’ self-reports reinforce this finding: intention alignment ratings—users’ perceived alignment between their activities and stated intentions—are higher under INA (4.44 vs. 4.23, $p<.001$). These results together indicate that users exhibited increased intentional activity with INA.
@@
Users also self-assessed their concentration over the previous week via the focused immersion scale in the post-survey. As shown in Figure 7(a), INA achieved higher scores than the baselines (3.74 vs. 3.34 for simple reminder, $p=.0449$; and 3.74 vs. 2.90 for logging only, $p=.0003$), providing further evidence that INA aided users’ focus. The focused immersion score further revealed significant differences across all three applications: even simple reminder outperformed logging only (3.34 vs. 2.90, $p=.0440$), suggesting that simply entering and being reminded of one’s intention already has a positive impact on focus, while INA, with its context-aware interventions, produced the largest improvement. In the interviews, all participants except one reported positive experiences with INA, noting improvements in work efficiency, such as “It [INA] helped me focus more quickly, so my work efficiency improved” (P15).
@@ 7.3. User Experience Survey
Figure 7. (a) Mean focused immersion scores, a self-reported measure of concentration taken during the weekly post-survey. (b) Mean user experience survey scores, computed by averaging responses within categories: Support , Message Effectiveness, Workflow Disruption. Error bars denote 95% confidence intervals across participants (user-averaged). ***, **, and *, indicate significance of p ¡ 0.001, p ¡ 0.01, p ¡ 0.05, respectively. ($\uparrow$) indicates that higher values are more favorable. Additional details are in Appendix D.3, including the scores for each question.:
@@
Figure 7(b) summarizes users’ overall experiences across three categories from the weekly post-survey. In the Support category, participants rated INA as significantly more effective than logging only in supporting intentional digital activities (3.75 vs. 2.75, $p<.001$). Compared to simple reminder, INA also received higher ratings for supporting intentional digital activities, although this difference did not reach statistical significance (3.75 vs. 3.56, $p=.267$). As shown in the Message Effectiveness category, users identified the context-aware, timely notifications of INA as a key factor underlying its benefits, as evidenced by higher scores for INA than for simple reminder (3.74 vs. 3.24, $p=.013$). In contrast, INA received the lowest score in the Workflow Disruption category (lower scores indicate more disruption) compared to simple reminder (3.09 vs. 3.91, $p=.005$) and logging only (3.09 vs. 3.91, $p=.005$). This aligns with the main concern raised in interviews, where participants noted that INA could be bothersome due to excessive interruptions.
@@ 7.4. Interview Findings
Figure 8. An illustration of INA’s safety challenges when faced with harmful intentions. (1) A user states a harmful or unsafe intention, such as ”Hacking into a hospital server.” (2) As the user performs actions aligned with this goal (e.g., visiting a medical records page of a hospital portal, watching a tutorial on phishing emails), INA provides positive reinforcement, encouraging the dangerous activity. (3) When the user’s behavior deviates to a safe but unrelated topic (e.g., watching an ethics lesson), the system perceives this as a distraction and delivers a notification, prompting the user to return to the original harmful intention. Such behavior occurs even though the underlying LLM is expected to refuse harmful requests, and shows the need to introduce additional safety measures.:
@@ 8.1. Summary of Findings
  • RQ2: What is the overall user experience with INA? In comparison to the other two applications, users perceived INA as more valuable for guiding their digital activities toward their intentions and experienced increased focus, but they also raised concerns about data privacy and excessive intervention (Sections 7.3 and 7.4).
@@ 8.2. Limitations and Future Directions
Reflected in positive interview and higher user experience survey ratings for support and effectiveness, participants accepted INA as a focus-restoring aid but also reported fatigue due to taxing aspects of its user experience such as Q&As being required for every session, feedback records lasting only for a day, and workflow disruption from frequent notifications. To prevent such fatigue, future work could adopt adaptive policies that personalize intervention frequency, timing, and intensity, with either user-facing controls for baseline preferences (e.g., message tone, timing) or automatic preference learning from interaction data. We believe that such personalization would substantially improve user acceptance and effectiveness of the system.
@@
Although we made strong efforts to anonymize the data and ensure careful handling, about half of the participants still expressed concerns regarding the treatment of sensitive information. These concerns suggest that technical protections alone may not be sufficient, and future work should explore building stronger privacy safeguards. For example, an on-device system that processes sensitive data locally, without transmitting it to external servers, could reduce the privacy risk.
Potential bias in the user study
During the orientation, participants were informed that one of the three applications included an AI agent, without specifying which role it occupied. Other minor differences and similarities, such as intention prompting and screenshot collections, were also introduced. To minimize potential bias and prevent participants from inferring features based on application names, the applications were color-coded from orientation through the study period. Nevertheless, we acknowledge that knowing one application involved an AI agent—as well as —and from that knowledge, perhaps assuming that the researchers preferred it— may have led participants to hold different attitudes towards and expectations for that version.
@@
Similar to the risks posed by AI agents (Andriushchenko et al., 2025; Lee et al., 2024a, 2025), INA also carries the risk of reinforcing unsafe intentions. In Figure 8, we illustrate an example case where INA provides positive reinforcement to a harmful intention “hacking into a hospital server”. Surprisingly, INA did not refuse to provide positive feedback to such harmful intentions, even though one might normally expect the underlying LLM to have the capability to reject them. We hypothesize that this behavior persists even when common refusal mechanisms are in place, since the LLM is tasked with predicting a user’s intentions rather than directly executing the harmful request. To mitigate such outcomes, we expect that integrating external guardrail models (e.g., WildGuard (Han et al., 2024)) can effectively detect harmful intentions and prevent the system from providing encouraging feedback.
@@ Acknowledgements.
This material is based upon work supported by the Google Cloud Research Credits program. This work has taken place in part in the Rewarding Lab at UT Austin. During this project, the Rewarding Lab has been supported by NSF (IIS-2402650), ONR (N00014-22-1-2204), ARO (W911NF-25-1-0254), Emerson, EA Ventures, UT Austin’s Good Systems grand challenge, and Open Philanthropy.
@@ A.1. Clarification
[…]
@@
[…]
@@ A.4. Feedback
[…]
A.5. System Architecture and Implementation
The INA system follows a client-server architecture, consisting of a client application running on the user’s local machine and a backend hosted on a cloud platform. We note that the implementation was aided by the LLM, especially when we code in Python.
Client Application
The client is a Python-based desktop application that is responsible for data transmission to and reception from the back-end.
  • Data Transmission: Transmits user data required for server-side analysis. This data includes (a) the user’s intention and related chat history, (b) periodically captured screen images, and (c) computer usage logs.
  • Interaction Processing: Receives the analysis results of the server, displays them in the UI, and transmits the user feedback back to the server. Key UI functions include presenting questions for intention clarification, displaying notifications after misalignment, and receiving user feedback.
Backend Infrastructure
The backend is developed in Python and deployed on a serverless platform (Cloud Run), which scales according to traffic. Its primary function is to synthesize data from the client and use Google’s Gemini 2.0- Flash model to infer the alignment level between the user’s current activity and their stated intention. It also performs the reflection process required for intention clarification and feedback processing via the LLM.
Data Infrastructure
Data is stored based on its role as follows:
  • Firestore: Stores structured data such as anonymized IDs, intentions, and feedback.
  • Cloud Storage: Stores unstructured data, such as de-identified screen images.
  • BigQuery: Utilized as a data warehouse for long-term pattern analysis.
@@ B.2. Evaluation Details
Figure 9. User interfaces for INA (top), simple reminder (middle), and logging only (bottom); INA and simple reminder conduct an Intention alignment rating survey at the end of the session.:
@@ C.3. Survey and Interview Instruments
Appendix D Further Analyses
D.1. Proportion of Off-task Time for Categorized Intention
@@ 9. Self-care and maintenance
Figure 10. Mean LLM-estimated off-task ratio ($\downarrow$) for productive intentions and non-productive intentions. Error bars denote 95% confidence intervals across participants (user-averaged). ***, **, and *, indicate significance of p ¡ 0.001, p ¡ 0.01, p ¡ 0.05, respectively. Arrows indicate whether higher ($\uparrow$) or lower ($\downarrow$) values are more favorable.:
Figure 11. Mean LLM-estimated off-task ratio shown by intention alignment rating (1–5). Error bars denote 95% confidence intervals across participants (user-averaged).:
D.3. User Experience Survey Results
@@
Figure 12. Results of 15-item user experience survey responses by application on a 5-point Likert scale. Means and standard deviations are shown on the right; (rev.) indicates reverse-coded items. ***, **, and *, indicate significance of p ¡ 0.001, p ¡ 0.01, p ¡ 0.05, respectively.:
D.4. Interview Coding Reliability Thematic Analysis
D.5. Open-ended Survey Coding Descriptive Analysis
@@
We applied descriptive coding to analyze the open-ended survey responses. This method assigns short labels to each response segment to summarize the experiences or perceptions reported by participants. The final codebook and results are presented in Table 7. The results illustrate that logging only provided unobtrusive logging with little support, simple reminder added reminders with mixed sufficiency, and INA delivered the strongest refocusing effects while introducing new challenges in precision and reliability.
Appendix E Privacy and Security Policies
Data De-identification
Access control
Data Encryption

### v3

@@ 7.2. Impact of INA on Intended Activity and User Focus
As shown in Figure 6, for the two applications that collect user intentions, INA not only reduces off-task time during task execution but also increases users’ reported alignment between their digital activity and their stated intention. The LLM-estimated off-task ratio [^6] Our technical evaluation in Section 5 provides evidence of its accuracy. is significantly lower with INA than with simple reminder (0.104 vs. 0.166, $\beta=-0.058$, 95% CI [0.082-0.082, 0.033-0.033], $p<.001$, $d=0.28$). Users’ self-reports reinforce this finding: intention alignment ratings—users’ perceived alignment between their activities and stated intentions—are higher under INA (4.44 vs. 4.23, $\beta=0.23$, 95% CI [$0.12$, $0.33$], $p<.001$, $d=0.26$). These results together indicate that users exhibited increased intentional activity with INA.
@@
Users also self-assessed their concentration over the previous week via the focused immersion scale in the post-survey. As shown in Figure 7(a), INA achieved higher scores than the baselines (3.74 vs. 3.34 for simple reminder, $\beta=0.40$, 95% CI [0.08-0.08, $0.88$], $p=.0449$, $d=0.62$; and 3.74 vs. 2.90 for logging only, $\beta=0.84$, 95% CI [$0.35$, $1.32$], $p=.0003$, $d=1.30$), providing further evidence that INA aided users’ focus. The focused immersion score further revealed significant differences across all three applications: even simple reminder outperformed logging only (3.34 vs. 2.90, $\beta=0.44$, 95% CI [0.05-0.05, $0.92$], $p=.0440$, $d=0.68$), suggesting that simply entering and being reminded of one’s intention already has a positive impact on focus, while INA, with its context-aware interventions, produced the largest improvement. In the interviews, all participants except one reported positive experiences with INA, noting improvements in work efficiency, such as “It [INA] helped me focus more quickly, so my work efficiency improved” (P15).
@@ 7.3. User Experience Survey
Figure 7. (a) Mean focused immersion scores, a self-reported measure of concentration taken during the weekly post-survey. (b) Mean user experience survey scores, computed by averaging responses within categories: Support , Message Effectiveness, Workflow Disruption. Error bars denote 95% confidence intervals across participants (user-averaged). ***, **, and *, indicate significance of p ¡ 0.001, p ¡ 0.01, p ¡ 0.05, respectively. ($\uparrow$) indicates that higher values are more favorable. Additional details are in Appendix D.3, including the scores for each question.
Bar charts comparing Logging Only, Simple Reminder, and INA on focused immersion and user experience survey scores.
@@
Figure 7(b) summarizes users’ overall experiences across three categories from the weekly post-survey. In the Support category, participants rated INA as significantly more effective than logging only in supporting intentional digital activities (3.75 vs. 2.75, $\beta=1.01$, 95% CI [$0.59$, $1.43$], $p<.001$, $d=1.81$). Compared to simple reminder, INA also received higher ratings for supporting intentional digital activities, although this difference did not reach statistical significance (3.75 vs. 3.56, $\beta=0.19$, 95% CI [0.23-0.23, $0.61$], $p=.267$, $d=0.34$). As shown in the Message Effectiveness category, users identified the context-aware, timely notifications of INA as a key factor underlying its benefits, as evidenced by higher scores for INA than for simple reminder (3.74 vs. 3.24, $\beta=0.50$, 95% CI [$0.12$, $0.88$], $p=.013$, $d=0.82$). In contrast, INA received the lowest score in the Workflow Disruption category (lower scores indicate more disruption) compared to simple reminder (3.09 vs. 3.91, $\beta=-0.82$, 95% CI [1.48-1.48, 0.15-0.15], $p=.0056$, $d=0.93$) and logging only (3.09 vs. 3.91, $\beta=-0.82$, 95% CI [1.48-1.48, 0.15-0.15], $p=.0056$, $d=0.93$). This aligns with the main concern raised in interviews, where participants noted that INA could be bothersome due to excessive interruptions.
@@ 8.1. Summary of Findings
  • RQ2: What is the overall user experience with INA? In comparison to the other two applications, users perceived INA as more valuable for guiding their digital activities toward their intentions and experienced increased focus, but they also raised concerns about data privacy and intrusiveness intervention (Sections 7.3 and 7.4).
@@
8.2. Implications of INA for Intentional Digital Living
Explaining INA with Dual-Systems theory
According to Dual-Systems theory (see Section 2.1), the intention-behavior gap in digital environments emerges when potent digital stimuli trigger impulsive System 1 reactions, displacing the goals of the resource-limited System 2 from working memory (Lyngs et al., 2019). From this perspective, INA augments System 2’s monitoring capacity in distraction-rich environments. By interpreting screen context in real-time, INA identifies and intervenes at the moment when an impulsive System 1 response is likely to override the user’s intentions. This effectively re-engages System 2 at critical moments where it fails, and attention is distracted. Our findings validate the importance of such timely intervention, consistent with prior research showing that early intervention is significantly more effective than suppressing responses after impulses have fully formed (Duckworth et al., 2016). Taken together, our work suggests that context-aware systems can meaningfully support intentional digital behavior without undermining user autonomy by preempting System 1 impulses and reinforcing System 2 control.
Respecting Diverse Intentions
The core mechanism of INA does not rely on pre-defined lists of productive applications. Instead, it utilizes an LLM to evaluate the semantic alignment between the user’s stated intention and their current behavior. This structural flexibility allows the system to treat non-productive intentions, such as “resting while watching YouTube”, with the same respect and support as productive goals, acknowledging the positive value of rest and play (Reinecke, 2009; Rieger et al., 2014; Valkenburg and Peter, 2007). Within this feature, conventional notions of distraction are inverted: when a user intends to take a full break, typical productive tools such as work messengers or document editors are detected as distractions that disrupt the intended rest. This approach goes beyond what existing DSCTs support, which have traditionally equated “on-task” states with productivity and treated rest as a behavior to be blocked (Lyngs et al., 2019; Roffarello and De Russis, 2023). Ultimately, INA illustrates how digital well-being tools can move beyond simple time-reduction strategies (Roffarello and De Russis, 2023; Vanden Abeele, 2020) to better support user autonomy, assisting whatever form of digital life the user defines.
Privacy and Ethical Challenges in AI-based DSCTs
Applying AI to DSCTs offers the promise of fostering healthier digital habits through personalized, context-aware interventions, but it also introduces substantial ethical and privacy challenges (Scibetta et al., 2025). Our study showed that some participants perceived the AI system not only as a tool but as a social entity. Although this perception may increase intervention acceptance, it risks encouraging emotional attachment or over-reliance on the system for self-regulation. This underscores the need for deeper reflection on the appropriate role of AI—whether as a tool, companion, or advisor—within the DSCT context.
Data privacy has also been consistently identified as a major concern in AI-based DSCTs (Scibetta et al., 2025). Despite our implementation of multiple safeguards, INA structurally depends on transmitting data to the cloud and trusting the policies of external AI API providers. Screen images introduce particular risks, as they may inadvertently contain sensitive information about interlocutors, colleagues, or organizations. This raises clear ethical and security issues if transmitted to external servers without explicit consent. Therefore, designing AI-based DSCTs should prioritize balancing strong privacy protections with effective, personalized well-being support. Privacy-by-design strategies (such as employing lightweight on-device LLMs (Qin et al., 2024) and pre-filtering sensitive on-screen content before transmission (Chen et al., 2025)) offer promising directions for mitigating these concerns.
@@ 8.3. Limitations and Future Directions
Reflected in positive interview and higher user experience survey ratings for support and effectiveness, participants accepted INA as a focus-restoring aid but also reported fatigue due to taxing aspects of its user experience such as Q&As being required for every session and workflow disruption from frequent notifications. To prevent such fatigue, future work could adopt dynamic elicitation processes and adaptive policies that personalize intervention frequency, timing, and intensity, with either user-facing controls for baseline preferences (e.g., message tone, timing) or automatic preference learning from interaction data. We believe that such personalization would substantially improve user acceptance and effectiveness of the system.
@@
Potential bias in the formative and user study
We acknowledge that the qualitative themes we identified may partially reflect biases associated with the demographic characteristics and digital usage habits of our participant group. The formative interviews were conducted with 11 participants recruited from a single university community, all of whom used computers for their work. This sampling strategy was appropriate for a qualitative design study aimed at identifying functional and interaction requirements for INA. However, the small sample size does not represent the broader population of computer users.
During the orientation, participants were informed that one of the three applications included an AI assistant, without specifying which role it occupied. Other minor differences and similarities, such as intention prompting and screenshot collections, were also introduced. To minimize potential bias and prevent participants from inferring features based on application names, the applications were color-coded from orientation through the study period. Nevertheless, we acknowledge that knowing one application involved an AI assistant (and potentially assuming that the researchers favored it) may have influenced participants’ attitudes toward and expectations for that version.
Limitations of IntentionBench
IntentionBench serves as a controlled benchmark that enables the systematic evaluation of INA and its components, helping us understand performance trade-offs and inform design decisions. However, it does not fully capture the complexity of natural off-task behaviors, such as multitasking across multiple windows, infinite scrolling, or internal distractions that are not visible on-screen. It also imposes limits on handling real-world ambiguity. For example, as illustrated in Figure 4, an activity strictly labeled as off-task (i.e., “watching e-sports”) could validly align with a user’s intention to “study HCI” in contexts such as analyzing interface design for e-sports. Expanding the dataset to include these natural off-task behaviors and ambiguous scenarios would be an interesting direction for future work and would enable more ecologically valid evaluations.
Diversity of Message Tones
In this study, we primarily adopted a supportive and encouraging tone to minimize perceptions of coercion. However, we did not explicitly isolate the effects of using different tonal styles, which may range from neutral or objective to playful or even authoritative. Prior work suggests that user responses can vary substantially depending on the emotional framing of an intervention (Saffaryazdi et al., 2025; Yun and Park, 2022). We believe that an important future direction is to systematically examine how such emotional framings interact with contextual timing to optimize both user acceptance and behavioral outcomes.
Figure 8. An illustration of INA’s safety challenges when faced with harmful intentions. (1) A user states a harmful or unsafe intention, such as ”Hacking into a hospital server.” (2) As the user performs actions aligned with this goal (e.g., visiting a medical records page of a hospital portal, watching a tutorial on phishing emails), INA provides positive reinforcement, encouraging the dangerous activity. (3) When the user’s behavior deviates to a safe but unrelated topic (e.g., watching an ethics lesson), the system perceives this as a distraction and delivers a notification, prompting the user to return to the original harmful intention. Such behavior occurs even though the underlying LLM is expected to refuse harmful requests, and shows the need to introduce additional safety measures.
A diagram showing an example of INA reinforcing a harmful intention.
@@
Similar to the risks posed by AI assistants (Andriushchenko et al., 2025; Lee et al., 2024, 2025), INA also carries the risk of reinforcing unsafe intentions. In Figure 8, we illustrate an example case where INA provides positive reinforcement to a harmful intention “hacking into a hospital server”. Surprisingly, INA did not refuse to provide positive feedback to such harmful intentions, even though one might normally expect the underlying LLM to have the capability to reject them. We hypothesize that this behavior persists even when common refusal mechanisms are in place, since the LLM is tasked with predicting a user’s intentions rather than directly executing the harmful request. To mitigate such outcomes, we expect that integrating external guardrail models (e.g., WildGuard (Han et al., 2024)) can effectively detect harmful intentions and prevent the system from providing encouraging feedback.
@@ Acknowledgements.
This work was supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (RS-2025-02304967, AI Star Fellowship (KAIST); RS-2019-II190075, Artificial Intelligence Graduate School Program (KAIST)). This research was partially supported by the National Science Foundation (NRT 2125858) and Good Systems, a UT Austin Grand Challenge for developing responsible AI technologies [^7] https://goodsystems.utexas.edu. This work has taken place in part in the Rewarding Lab at UT Austin. During this project, the Rewarding Lab has been supported by NSF (IIS-2402650), ONR (N00014-22-1-2204), ARO (W911NF-25-1-0254), Emerson, EA Ventures, and Open Philanthropy. This award is with support from Google.org and the Google Cloud Research Credits program for the Gemma Academic Program.
@@ A.1. Clarification
[Output] Only provide your question in a single sentence (at most 10 words).
@@
[Output] Return only the JSON object, with keys ”1”–”10”; no additional text.
@@ A.4. Feedback
Only return the JSON object; do not include any additional text.
A.5. System Architecture and Implementation Details
The INA system follows a client-server architecture, comprising a client application on the user’s local machine and a backend hosted on a cloud platform. The implementation was aided by LLMs (e.g., Python code generation).
Client Application and Libraries
The client is a Python-based native macOS application. To address the technical requirements for context-awareness, we utilized specific libraries tailored for macOS integration. Table 3 summarizes the key libraries.
Table 3. Key libraries used in the INA client application.
Category Library Purpose
UI Framework PyQt6 Main application window and system tray
OS Integration rumps macOS Menu bar application framework
Screen/Session pyobjc Screen capture and session lock detection
Image Proc. Pillow Screenshot compression and resizing
Notifications desktop-notifier Native macOS notifications with actions
LLM Client google-genai Direct communication with Gemini API
Data Collection and Context Detection
The client captures user activity logs at specific intervals to infer context. Table 4 details the collected log types.
Table 4. Data types and collection methods employed by the client.
Data Type Method Int. Note
Screenshot PyQt6 2s Compressed JPEG (Quality=85)
Active App AppleScript 2s Name of the foreground app
Browser URL AppleScript 2s URL of the active tab
App Change Delta Check Evt. Triggered on app/domain switch
Screen Lock Quartz API 1s Auto-pauses session if locked
URL Detection Mechanism
As web browsing is a significant part of digital activity, precise URL detection is crucial. INA utilizes AppleScript via the osascript bridge to fetch the URL from the active tab of supported browsers. For example, the command for Google Chrome is:
tell application "Google Chrome" to get URL of active tab of front window
The extracted URL is parsed to identify the domain. Changes in the domain are treated as context switches, prompting the LLM to re-evaluate the user’s alignment.
Backend and Data Infrastructure
The backend is built with FastAPI on Google Cloud Run. It synthesizes the client data and utilizes the Gemini 2.0-Flash model (Temperature=0.1) to infer the alignment between the user’s current activity and their stated intention. Data is stored securely based on its type:
  • Firestore: Structured data (anonymized IDs, intentions, feedback, logs).
  • Cloud Storage: Unstructured data (de-identified screen images).
  • BigQuery: Data warehouse for long-term pattern analysis.
  • Local Storage: JSON-based session history and reflection logs are also cached locally on the client for performance.
@@ C.1. Pre-survey
Figure 9. User interfaces for INA (top), simple reminder (middle), and logging only (bottom); INA and simple reminder conduct an Intention alignment rating survey at the end of the session.
Interfaces screenshot of INA, simple reminder, and logging only systems.
@@ 9. Self-care and maintenance
Appendix D Further Analyses
D.1. Proportion of Off-task Time for Categorized Intention
D.3. User Experience Survey Results
D.4. Interview Coding Reliability Thematic Analysis
D.5. Open-ended Survey Coding Descriptive Analysis
We applied descriptive coding to analyze the open-ended survey responses. This method assigns short labels to each response segment to summarize the experiences or perceptions reported by participants. The final codebook and results are presented in Table 9. The results illustrate that logging only provided unobtrusive logging with little support, simple reminder chadded reminders with mixed sufficiency, and INA delivered the strongest refocusing effects while introducing new challenges in precision and reliability.
Appendix E Privacy and Security Policies
Data De-identification
Access control
Data Encryption
@@ Appendix E Privacy and Security Policies
Figure 10. Mean LLM-estimated off-task ratio ($\downarrow$) for productive intentions and non-productive intentions. Error bars denote 95% confidence intervals across participants (user-averaged). ***, **, and *, indicate significance of p ¡ 0.001, p ¡ 0.01, p ¡ 0.05, respectively. Arrows indicate whether higher ($\uparrow$) or lower ($\downarrow$) values are more favorable.
Bar charts comparing INA and simple reminder on off-task ratios for productive and non-productive intentions.
Figure 11. Mean LLM-estimated off-task ratio shown by intention alignment rating (1–5). Error bars denote 95% confidence intervals across participants (user-averaged).
A bar chart depicting the relationship between intention alignment ratings (1–5) and LLM-estimated off-task ratios.
@@
Figure 12. Results of 15-item user experience survey responses by application on a 5-point Likert scale. Means and standard deviations are shown on the right; (rev.) indicates reverse-coded items. ***, **, and *, indicate significance of p ¡ 0.001, p ¡ 0.01, p ¡ 0.05, respectively.
Bar charts showing user experience survey responses across INA, simple reminder, and logging only system.